# A Unified Framework for Input Feature Attribution Analysis

### v1

@@ Abstract
Explaining the decision-making process of machine learning models is crucial for ensuring their reliability and fairness. One popular explanation form highlights key input features, such as i) tokens (e.g., Shapley Values and Integrated Gradients), ii) interactions between tokens (e.g., Bivariate Shapley and Attention-based methods), or iii) interactions between spans of the input (e.g., Louvain Span Interactions). However, these explanation types have only been studied in isolation, making it difficult to judge their respective applicability. To bridge this gap, we propose a unified framework that facilitates a direct comparison between highlight and interactive explanations comprised of four diagnostic properties [^1] We make the code for the framework available at https://github.com/copenlu/A-unified-framework-for-highlight-explanations .. Through extensive analysis across these three types of input feature explanations–each utilizing three different explanation techniques–across two datasets and two models, we reveal that each explanation type excels in terms of different diagnostic properties. In our experiments, highlight explanations are the most faithful to a model’s prediction, and interactive explanations provide better utility for learning to simulate a model’s predictions. These insights further highlight the need for future research to develop combined methods that enhance all diagnostic properties.
@@ 1 Introduction
Figure 1: An example of the three types of Input Feature Explanations on an instance from the SNLI dataset, with their two most important pieces of explanation (token, token tuple, span tuple, correspondingly).
@@
Input feature explanations reveal how a model makes decisions based on the input for which the prediction is made. Among these, Token Explanations (TokenEx; e.g., Shapley Values Lundberg and Lee (2017) and Integrated Gradients (Sundararajan et al., 2017)) are widely used to assign importance scores to specific segments of the input, such as tokens in natural language processing tasks. However, complex reasoning tasks usually require reasoning across multiple pieces of text, e.g., Fact Checking is performed given a claim and evidence, Natural Language Inference is performed given a premise and a hypothesis, where Token Explanations can be insufficient to present the relations employed between the different parts of the input. To this end, Token Interactive Explanations (TokenIntEx; e.g., Bivariate Shapley Masoomi et al. (2023), and Layer-wise Attention Attribution (Ye et al., 2021)) provide importance scores for interactions between two tokens of the input. Further, to enhance the semantic coherence of these explanations, Span Interactive Explanations (SpanIntEx; e.g., Louvain community detection (Ray Choudhury et al., 2023) model interactions across tuples of input spans. An example showcasing these three types of input feature explanations is given in Figure 1.
@@
It is crucial to develop rigorous and comprehensive evaluation frameworks to ensure the principled selection of the most suitable explanation in a practical application Yu et al. (2024) and systematic progress in the development of different types of explanations Atanasova et al. (2022); Jolly et al. (2022). However, these input feature explanations have typically been studied in isolation, with few exceptions, where explanation methods of the same type are compared Atanasova et al. (2020); DeYoung et al. (2020); Janizek et al. (2021); Ray Choudhury et al. (2023). Moreover, evaluations of interactive explanations have been restricted to one property. To address these gaps, we propose a unified framework that facilitates a direct comparison between highlight and interactive input feature explanations on a suite of four diagnostic properties.
@@
Unified Evaluation Framework. Our unified evaluation framework consists of four essential diagnostic properties – Faithfulness, Agreement with Human Annotation, Simulatability, and Complexity. They are the most widely used for Token explanations and include the only property used for interactive explanations – Faithfulness. Faithfulness (Chen et al., 2020, 2021; Ray Choudhury et al., 2023) measures the extent to which explanations accurately reflect the actual reasons used by the model in its predictions. Agreement with Human Annotation (Atanasova et al., 2020) evaluates whether explanations exhibit an inductive bias akin to human reasoning, potentially enhancing their plausibility to end users. Simulatablity (Pruthi et al., 2022) estimates whether the explanations are useful to an agent for replicating the model’s decisions. Finally, Complexity (Bhatt et al., 2021) evaluates whether the explanations are easy to comprehend by presenting distinctive key features rather than a uniform distribution of importance scores. Crucially, we propose extensions of these properties such that they can be applied to different types of input feature explanations – token (TokenEx), two tokens (TokenIntEx), or more than two tokens (SpanIntEx).
@@
Characterising Input Feature Explanations. We conduct an extensive analysis across two different tasks and language models, and three different explanation techniques for each input feature explanation type. Our findings indicate that TokenEx explanations exhibit greater faithfulness than SpanIntEx and TokenIntEx explanations. Additionally, SpanIntEx and TokenIntEx explanations align more closely with human annotations at the token level compared to TokenEx explanations. Moreover, SpanIntEx explanations demonstrate higher interaction overlap with human annotations than TokenIntEx explanations. Further, SpanIntEx explanations are found to be most beneficial for agent simulatability. Finally, our results suggest that SpanIntEx and TokenEx explanations tend to be easier to comprehend.
Overall, our extensive analysis reveals the varied strengths of each explanation type across different diagnostic properties. These insights underscore the importance of developing integrated methods that combine the different types of explanations to enhance all diagnostic properties in future research.
2 A Unified Evaluation Framework for Highlight Explanations
@@
To systematically compare feature attribution explanations of different types, we present a unified framework comprised of four widely employed diagnostic properties: Faithfulness, Agreement with Human Annotations, Simulatability, and Complexity. This section formally introduces these properties and outlines our proposed extensions to allow for their application across different input feature explanation types.
@@ 2.1 Preliminaries
We start with a dataset $D$, and a model $M$ fine-tuned on $D$. An instance $x\in D$ is comprised of two parts, e.g., a claim and an evidence, the first consisting of $m$ tokens, and the second – of $n$ tokens. We apply an explanation attribution method $A_{E}$ of type $E\in\{\texttt{TokenEx},\texttt{TokenIntEx},\texttt{SpanIntEx}\}$ to $M$, and each instance $x\in D$: $A_{E}(M,x)=\{e^{x}_{k},a^{x}_{k}|k\in[0,s-1]\}$, where $e^{x}_{k}$ is a span/a pair of spans from the input and $a^{x}_{k}$ denotes its importance score. $k$ is the importance ranking of the corresponding piece of explanation. $s$ is an upper limit for the number of most important pieces of explanation for an instance, such that: $s\in[1,m+n]$ for TokenEx, $s\in[1,m\cdot n]$ for TokenIntEx, and for SpanIntEx $s$ varies for each instance with $s\in[1,f],f<m!\cdot n!$. Depending on the explanation type $E$, $e^{x}_{k}$ can consist of: one token $x_{i}$ for TokenEx, $i\in[0,m+n-1]$, one token pair $(x_{p},x_{q})$ for TokenIntEx, where $p\in[0,m-1]$ and $q\in[m,m+n-1]$, one span pair $((x_{s},\dots,x_{s+l_{1}}),(x_{t},\dots,x_{t+l_{2}}))$ for SpanIntEx, where $s,s+l_{1}\in[0,m-1]$ and $t,t+l_{2}\in[m,m+n-1]$.
@@
$T_{E,M}(x,s)=\{x_{i}|x_{i}\in{e}_{k}^{x},k\in[0,s-1]\}$ (1)
When computing the different diagnostic properties, previous works have adopted a range of thresholds $s$. Moreover, as noted above, the maximum possible value for $s$ can vary across different types of input feature explanations. Additionally, the number of tokens included in the top-$k$ important explanations can differ substantially among explanation types: in top-1 we can have: in a top-1 setting, the explanations consist of 1 token for TokenIntEx, 2 tokens for TokenIntEx, and more than 2 tokens for SpanIntEx Ray Choudhury et al. (2023); DeYoung et al. (2020). This variability makes it difficult to compare results across different explanation types. To address this problem, we propose extensions for each of the studied diagnostic properties that result in unified diagnostic properties which can be used for a direct comparison of the different types of input feature explanations.
@@ 2.2 Faithfulness
Faithfulness DeYoung et al. (2020) assesses whether explanations accurately reflect the model’s decision-making process. It involves two aspects – Comprehensiveness and Sufficiency – measured as the number of the model’s prediction changes after keeping (SP) or omitting (CP) $k$ of the most important portions of the input, respectively:
$CP(x,E,k)=\left\{\begin{array}[]{ll}0,&\text{if }f(x-T_{E,M}(x,k))=f(x)\\ 1,&\text{otherwise }\end{array}\right\}$ (2)
$SP(X,E,k)=\left\{\begin{array}[]{ll}1,&\text{if }f(T_{E,M}(x,k))=f(x)\\ 0,&\text{otherwise }\end{array}\right\}$ (3)
@@
To take different thresholds $k$, we average over $k\in[0,s=m+n-1]$. We then also average the results across instances within $D$ to compute the final Comprehensiveness/Sufficiency score:
$Comp_{E}/Suff_{E}=\frac{\sum_{x\in{D}}^{|D|}\sum_{k=0}^{s}CP/SP(x,E,k)}{|D|}$ (4)
@@
Unified Faithfulness. To facilitate the comparison of faithfulness scores, we introduce a dynamic threshold $\theta_{x,k}$. This threshold represents the number of unique tokens used for a perturbation on input $x$, same across all explanation types. Since top-$k$ explanations from SpanIntEx typically contain more tokens than TokenEx or TokenIntEx, the number of tokens in top-$k$ SpanIntEx explanations sets the number of unique tokens used for perturbations across all types of input feature explanations:
$\theta_{x,k}=\left|T_{\texttt{SpanIntEx},M}(x,k)\right|$ (5)
@@
Thus, $\theta_{x,k}$ becomes a dynamic threshold that adapts based on each instance’s top-$k$ SpanIntEx explanation tokens. We then adjust the number of top-$k$ explanations selected from TokenEx and TokenIntEx, $k_{\texttt{TokenEx}}$ and $k_{\texttt{TokenIntEx}}$, correspondingly, to result in the same number of perturbed tokens $\theta_{x,k}$:
$\left|T_{\texttt{TokenEx},M}(x,k_{\texttt{TokenEx}}(x))\right|=\theta_{x,k}$ (6)
$\left|T_{\texttt{TokenIntEx},M}(x,k_{\texttt{TokenIntEx}}(x))\right|=\theta_{x,k}$ (7)
@@
Furthermore, a Random baseline is established, where the tokens for perturbation are selected randomly so that it matches the average $\theta_{x,k}$ across $D$.
2.3 Agreement with Human Annotation
@@
Agreement with Human Annotation has been used to assess the overlap between generated and human-annotated TokenEx explanations Atanasova et al. (2020), which can indicate the plausibility of the generated explanations to end users. For $E=\texttt{TokenEx}$, the measure is computed by calculating a precision score for the top-$k$ most important explanations compared to the gold Human Annotations.
For $E=\texttt{TokenEx}$, $s$ are explanations extracted from instance $x$, where $s=m+n$, $s$ thresholds are set corresponding to each explanation’s attribution score, forming the threshold list $\omega_{E}(x)$:
$\omega_{E}(x)=[a^{x}_{0},...,a^{x}_{s}]$ (8)
@@
By selecting explanations with higher attribution scores than each threshold in $\omega_{E}(x)$, $s$ targeted explanation sets are obtained, where $C_{E}(x,i)$ represents the set for the $i^{th}$ threshold:
$C_{E}(x,i)=\{e^{x}_{j}:a^{x}_{j}<=a^{x}_{i}\}$ (9)
@@
Comparing these sets with the golden explanation set $e^{G}$, $s$ precision-recall pairs can be derived. Average Precision (AP) is then obtained by weighting the precision value with the corresponding recall increase:
$P_{i}/R_{i}(x,e^{G},E)=Pre/Rec(C_{E}(x,i),e^{G})$ (10)
$AP_{E}(x,e^{G})=\sum_{i=0}^{s}(R_{i}-R_{i-1})*P_{i}$ (11)
Finally, Mean AP (MAP) is calculated by averaging AP across the instances in the dataset $D$:
$MAP_{E}=\frac{\sum_{x\in D}^{|D|}AP_{E}(x,e^{G})}{|D|}$ (12)
@@
Unified Agreement with Human Annotation Measure. For a fair comparison between the different types of explanations, the thresholds $\omega_{E}(x)$ for including the same number of tokens across the different types of explanations follows the procedure set in the Unified Faithfulness property (§2.2). In particular, we start with the thresholds set for SpanIntEx, and adjust the thresholds for TokenEx/TokenIntEx to reach the same number corresponding number of tokens. Furthermore, we measure agreement with human annotations at the interaction level for the gold SpanIntEx/TokenIntEx explanations and at the token level for gold TokenEx explanations.
Interaction-level Agreement with Human Annotation.
@@
For a fair comparison between TokenIntEx and SpanIntEx explanations, we adapt $MAP_{\texttt{TokenEx}}$ to the interaction level.
We compute the mean average precision (Eq. 12) w.r.t. the human-annotated TokenIntEx/SpanIntEx sets.
Token-level Agreement with Human Annotation.
@@
For a fair comparison between TokenEx and TokenIntEx/SpanIntEx explanations, we extract tokens from TokenIntEx/SpanIntEx explanations and compare them with tokens from golden TokenIntEx/SpanIntEx explanations. The threshold lists $\omega_{\texttt{TokenIntEx}/\texttt{SpanIntEx}}(x)$ are used, but the targeted sets $C_{E_{token}}(x,i)$ contain tokens extracted from TokenIntEx/SpanIntEx, correspondingly.
The golden set $S_{E_{token}}(x)$ aggregates tokens from golden TokenIntEx/SpanIntEx explanations. The token-level agreement scores $MAP_{\texttt{TokenIntEx}_{token}}/MAP_{\texttt{SpanIntEx}_{token}}$ accordingly, following the procedure set for $MAP_{\texttt{TokenEx}}$ (Eq. 12).
@@
We also set a Random baseline, where the number of randomly selected span pairs, token pairs, or tokens for each instance matches the average number of SpanIntEx explanations per instance across all attribution methods.

### v2

# Evaluating Input Feature Explanations through a Unified Diagnostic Evaluation Framework

@@ Abstract
Explaining the decision-making process of machine learning models is crucial for ensuring their reliability and transparency for end users. One popular explanation form highlights key input features, such as i) tokens (e.g., Shapley Values and Integrated Gradients), ii) interactions between tokens (e.g., Bivariate Shapley and Attention-based methods), or iii) interactions between spans of the input (e.g., Louvain Span Interactions). However, these explanation types have only been studied in isolation, making it difficult to judge their respective applicability. To bridge this gap, we develop a unified framework that facilitates an automated and direct comparison between highlight and interactive explanations comprised of four diagnostic properties [^1] Our code can be found at https://github.com/copenlu/A-unified-framework-for-input-feature-explanations. We conduct an extensive analysis across these three types of input feature explanations–each utilizing three different explanation techniques–across two datasets and two models, and reveal that each explanation has distinct strengths across the different diagnostic properties. Nevertheless, interactive span explanations outperform other types of input feature explanations across most diagnostic properties. Despite being relatively understudied, our analysis underscores the need for further research to improve methods generating these explanation types. Additionally, integrating them with other explanation types that perform better in certain characteristics could further enhance their overall effectiveness.
@@ 1 Introduction
Figure 1: An example of the three types of Input Feature Explanations on an instance from the SNLI dataset, with their two most important pieces of explanation (token for TokenEx, token tuple for TokenIntEx, span tuple for SpanIntEx, correspondingly).
@@
Input feature explanations reveal how a model makes decisions based on a specific input. Among these, the most widely used explanation type is Token Explanations (TokenEx), which for Natural language Understanding tasks provide importance scores for the tokens of the input, using methods such as Shapley Values Lundberg and Lee (2017) and Integrated Gradients (Sundararajan et al., 2017). However, for complex reasoning tasks that require reasoning across multiple pieces of text, e.g., Fact Checking is performed given a claim and evidence, Natural Language Inference is performed given a premise and a hypothesis, TokenEx can be insufficient to capture the relations employed between the different parts of the input. To this end, Token Interactive Explanations (TokenIntEx) are proposed as another explanation type that provides importance scores for interactions between two tokens of the input with methods such as Bivariate Shapley Masoomi et al. (2022) and Layer-wise Attention Attribution (Ye et al., 2021). Further, to enhance the semantic coherence of these interactive explanations, Span Interactive Explanations (SpanIntEx) is an explanation type that provides importance scores for interactions across tuples of spans, e.g., generated by Louvain community detection (Ray Choudhury et al., 2023). Figure 1 showcases these three types of input feature explanations.
@@
It is crucial to develop automated, rigorous and comprehensive evaluation frameworks to ensure the principled selection of the most suitable explanation in a practical application Yu et al. (2024) and systematic progress in the development of different types of explanations Atanasova et al. (2022); Jolly et al. (2022). However, these three types of input feature explanations have typically been studied in isolation, where explanation methods of the same type are compared Atanasova et al. (2020a); DeYoung et al. (2020); Janizek et al. (2021); Ray Choudhury et al. (2023). Moreover, evaluations of interactive explanations have been restricted to one property. To address these gaps, we develop a unified framework that facilitates an automated and direct comparison between highlight and interactive input feature explanations on a suite of four diagnostic properties. The framework allows for a rapid, scalable comparison across explainability methods, essential for the growing number of new techniques. Using the framework, we then perform an extensive empirical analysis of the properties of existing explanation methods across all three explanation types.
@@
Unified Evaluation Framework. Our unified evaluation framework consists of four essential diagnostic properties – Faithfulness, Agreement with Human Annotation, Simulatability, and Complexity. They are the most widely used for TokenEx Nauta et al. (2023) and include the only property used for interactive explanations – Faithfulness. Faithfulness (Chen et al., 2020, 2021; Ray Choudhury et al., 2023) measures the extent to which explanations accurately reflect the reasons used by the model in its predictions. Agreement with Human Annotation (Atanasova et al., 2020a) evaluates whether explanations exhibit an inductive bias akin to human reasoning, potentially enhancing their plausibility to end users. Simulatablity (Pruthi et al., 2022) estimates whether the explanations are useful to an agent for replicating the model’s decisions. Finally, Complexity (Bhatt et al., 2021) evaluates whether the explanations are easy to comprehend. In the unified evaluation framework, we further extend the four properties to facilitate their application and comparison across all three explanation types (§2).
@@
Extensive Empirical Analysis of Input Feature Explanations. We conduct an extensive analysis across two different textual tasks, two language models, and three explanation techniques for each input feature explanation type. Our findings indicate that TokenEx exhibit greater Comprehensiveness, and SpanIntEx – Sufficiency. Additionally, SpanIntEx and TokenIntEx align more closely with human annotations at the token level than TokenEx. Moreover, SpanIntEx demonstrate the highest interaction overlap with human annotations. Further, SpanIntEx are found to be most beneficial for Simulatability. Finally, our results suggest that SpanIntEx and TokenEx are easier to comprehend.
Overall, our analysis highlights the strengths of each explanation type across various diagnostic properties, with SpanIntEx generally outperforming others on most measures. However, we observe a trade-off between Comprehensiveness and plausibility, particularly with SpanIntEx, underscoring the need for methods that enhance both. Future research could explore integrated approaches that combine explanation types to optimize performance across all diagnostic properties.
2 A Unified Automated Evaluation Framework for Highlight Explanations
@@
According to the established taxonomy of evaluation approaches for explainability methods, evaluations can be functionally grounded, human-grounded, or application-grounded (Doshi-Velez and Kim, 2017). Our work focuses on automated functionally-grounded evaluation, which serves as a foundation for the other evaluation types as it enables the rapid, scalable comparison across explainability methods, essential for the growing number of new techniques. To this end, we present a unified framework comprised of four widely employed diagnostic properties: Faithfulness, Agreement with Human Annotations, Simulatability, and Complexity. This section formally introduces them and outlines the extensions that allow for their application across different input feature explanation types.
@@ 2.1 Preliminaries
We start with a dataset $D$, and a model $M$ fine-tuned on $D$. An instance $x\in D$ comprises two parts, e.g., a claim and an evidence, the first consisting of $m$ tokens, and the second – of $n$ tokens. We apply an explanation attribution method $A_{E}$ of type $E\in\{\texttt{TokenEx},\texttt{TokenIntEx},\texttt{SpanIntEx}\}$ to $M$, and each $x\in D$: $A_{E}(M,x)=\{e^{x}_{k},a^{x}_{k}|k\in[0,s-1]\}$, where $e^{x}_{k}$ is a token/pair of tokens/pair of token spans from the input and $a^{x}_{k}$ denotes its importance score. $k$ is the importance ranking of the corresponding piece of explanation. $s$ is an upper limit for the number of most important pieces of explanation for an instance, such that: $s\in[1,m+n]$ for TokenEx, $s\in[1,m\cdot n]$ for TokenIntEx, and for SpanIntEx $s$ varies for each instance with $s\in[1,f],f<m!\cdot n!$. Depending on the explanation type $E$, $e^{x}_{k}$ can consist of: one token $x_{i}$ for TokenEx, $i\in[0,m+n-1]$, one token pair $(x_{p},x_{q})$ for TokenIntEx, where $p\in[0,m-1]$ and $q\in[m,m+n-1]$, one span pair $((x_{s},\dots,x_{s+l_{1}}),(x_{t},\dots,x_{t+l_{2}}))$ for SpanIntEx, where $s,s+l_{1}\in[0,m-1]$ and $t,t+l_{2}\in[m,m+n-1]$.
@@
$T_{A_{E},M}(x,s)=\{x_{i}|x_{i}\in{e}_{k}^{x},k\in[0,s-1]\}$ (1)
However, as noted above, the upper bound for $s$ can vary across input feature explanations. Additionally, the number of tokens included in the top-$k$ important explanations can differ substantially among explanation types – in top-1 we can have: 1 token for TokenIntEx, 2 tokens for TokenIntEx, and more than 2 tokens for SpanIntEx. This variability makes it difficult to compare results across different explanation types. To this end, we propose extensions for each of the studied diagnostic properties that result in unified diagnostic properties applicable for a direct comparison of the different types of input feature explanations.
@@ 2.2 Faithfulness
Faithfulness DeYoung et al. (2020) assesses whether explanations accurately reflect the model’s decision-making process. It involves two aspects – Sufficiency and Comprehensiveness – measured as the number of the model’s prediction changes after keeping (SP) or omitting (CP) $k$ most important portions of the input (see omission details in §A):
$CP(x,A_{E},k)=\left\{\begin{array}[]{ll}0,&\text{if }f(x-T_{A_{E},M}(x,k))=f(x)\\ 1,&\text{otherwise }\end{array}\right\}$ (2)
$SP(X,A_{E},k)=\left\{\begin{array}[]{ll}1,&\text{if }f(T_{A_{E},M}(x,k))=f(x)\\ 0,&\text{otherwise }\end{array}\right\}$ (3)
@@
To take different thresholds $k$, we average over $k\in[0,s=m+n-1]$. We then also average the results across instances within $D$ to compute the final Comprehensiveness and Sufficiency scores:
$Comp_{A_{E}}=\frac{\sum_{x\in{D}}^{|D|}\sum_{k=0}^{s}CP(x,A_{E},k)}{s*|D|}$ (4)
$Suff_{A_{E}}=\frac{\sum_{x\in{D}}^{|D|}\sum_{k=0}^{s}SP(x,A_{E},k)}{s*|D|}$ (5)
@@
Unified Faithfulness. To facilitate the comparison of faithfulness scores, we introduce a dynamic threshold $\theta_{x,k}$. It represents the number of unique tokens used for a perturbation on $x$, same across explanation methods of all explanation types. Since top-$k$ explanations for $A_{E}$ of type SpanIntEx typically contain more tokens than for TokenEx or TokenIntEx, we set $\theta_{x,k}$ across the explanation methods of all types to:
$\theta_{x,k}=\left|T_{A_{\texttt{SpanIntEx}},M}(x,k)\right|$ (6)
@@
Thus, $\theta_{x,k}$ becomes a dynamic threshold that adapts based on each instance’s top-$k$ explanation tokens from $A_{\texttt{SpanIntEx}}$. We then adjust the number of top-$k$ explanations selected from $A_{\texttt{TokenEx}}$ and $A_{\texttt{TokenIntEx}}$, $k_{A_{\texttt{TokenEx}}}$ and $k_{A_{\texttt{TokenIntEx}}}$, correspondingly, to result in the same number of perturbed tokens $\theta_{x,k}$:
$\left|T_{A_{\texttt{TokenEx}},M}(x,k_{A_{\texttt{TokenEx}}}(x))\right|=\theta_{x,k}$ (7)
$\left|T_{A_{\texttt{TokenIntEx}},M}(x,k_{A_{\texttt{TokenIntEx}}}(x))\right|=\theta_{x,k}$ (8)
@@
Furthermore, a Random baseline is established, where the tokens for perturbation are selected randomly to match the average $\theta_{x,k}$ across $D$.
2.3 Agreement with Human Annotations
@@
Agreement with Human Annotations has been used to assess the overlap between generated and human-annotated explanations, which can indicate the plausibility of the generated explanations to end users. For $E=\texttt{TokenEx}$, the measure is computed by calculating a precision score for the top-$k$ most important explanations compared to the gold human annotations Atanasova et al. (2020a).
@@
For $E=\texttt{TokenEx}$, $a_{i}^{x},i\in[1,m+n]$ is the attribution score of $i$th most important explanation for $x$. $s=m+n$ is the number of explanations extracted from $x$. Corresponding to each explanation’s attribution score, $s$ thresholds are set, forming the threshold list $\omega_{A_{E}}(x)=[a^{x}_{0},...,a^{x}_{s}]$. By selecting explanations with higher attribution scores than each threshold in $\omega_{A_{E}}(x)$, $s$ targeted explanation sets are obtained, where $C_{A_{E}}(x,i)\{e^{x}_{j}:a^{x}_{j}<=a^{x}_{i}\}$ represents the set for the $i^{th}$ threshold, $a^{x}_{j}$ is the attribution score of token $e^{x}_{j}$ for $E=\texttt{TokenEx}$.
@@
Comparing these sets with the golden explanation set $e^{G}$, $s$ precision-recall pairs $P_{i}/R_{i}(x,e^{G},A_{E})$ can be derived. Average Precision (AP) is then obtained by weighting $P_{i}$ with the corresponding $R_{i}$ increase:
$P_{i}/R_{i}(x,e^{G},A_{E})=Pre/Rec(C_{A_{E}}(x,i),e^{G})$ (9)
$AP_{A_{E}}(x,e^{G})=\sum_{i=0}^{s}(R_{i}-R_{i-1})*P_{i}$ (10)
Mean AP (MAP) is calculated for all $x\!\in\!D$:
$MAP_{A_{E}}=\frac{\sum_{x\in D}^{|D|}AP_{A_{E}}(x,e^{G})}{|D|}$ (11)
@@
Unified Agreement with Human Annotation Measure. For a fair comparison between the different types of explanations, the thresholds $\omega_{A_{E}}(x)$ for including the same number of tokens across the explanation methods follows the procedure set for the Unified Faithfulness (§2.2).
Furthermore, we measure Agreement with Human Annotations at the interaction level for the gold SpanIntEx/TokenIntEx explanations and at the token level for gold TokenEx explanations.
@@
Interaction-level Agreement with Human Annotations. For a fair comparison between TokenIntEx and SpanIntEx methods, we adapt $MAP_{\texttt{TokenEx}}$ to the interaction level. Specifically, we compute the mean average precision (Eq. 11) w.r.t. the human-annotated TokenIntEx/SpanIntEx sets.
@@
Token-level Agreement with Human Annotations. For a fair comparison between TokenEx and TokenIntEx/SpanIntEx methods, we extract tokens from TokenIntEx/SpanIntEx and compare them with tokens from golden TokenIntEx/SpanIntEx sets. To compute MAP at token level, we follow the similar procedure set for $E=\texttt{TokenEx}$ (Eq. 11) with threshold lists $\omega_{A_{\texttt{TokenIntEx}/\texttt{SpanIntEx}}}(x)$, but the targeted sets $C_{A_{E_{token}}}(x,i)$ contain tokens extracted from TokenIntEx/SpanIntEx methods. The golden set $S_{A_{E_{token}}}(x)$ aggregates tokens from golden TokenIntEx/SpanIntEx sets.
@@
We also set a Random baseline, where the number of randomly selected span pairs, token pairs, or tokens for each instance matches the average number of tokens per instance extracted with a SpanIntEx explanation method.